OpenStudy (erniellyd28):

is 3 a solution 4y minus 8 equals 8y minus 2 - 4

OpenStudy (anonymous):

are you asking if 3/4 is a solution to this equation ? Lets find out.... 4y - 8 = 8y - 2 -8 + 2 = 8y - 4y - 6 = 4y -6/4 = y - 3/2 No, 3/4 is not s solution

OpenStudy (erniellyd28):

thank you so very much for all your help but if its ok may i please ask for your help on one more problem

OpenStudy (anonymous):

you can ask....I can't make no promises on whether I can do it

OpenStudy (erniellyd28):

is 2 a solution of 3m plus 7 equals 9m minus 3 - 3

OpenStudy (anonymous):

3m + 7 = 9m - 3 3m - 9m = -3 - 7 - 6m = - 10 m = -10/-6 m = 5/3 2/3 is not a solution

OpenStudy (erniellyd28):

thank you so very much for your help i really appreciate it